# Rho-1: Not All Tokens Are What You Need

### v1

@@ 4 Related Works
The objective of optimizing pre-training corpora is to maximize the performance and efficiency of language model training by improving the quality and scale of the pretrain data mixture. This typically includes data collecting through crawling (Raffel et al., 2020) or synthesis (Polu and Sutskever, 2020; Gunasekar et al., 2023), de-duplication (Lee et al., 2021; Kandpal et al., 2022; Tirumala et al., 2023), filtering and selection (Albalak et al., 2024), as well as data composition (Xie et al., 2024) and curriculum (Chen et al., 2024; MA et al., 2024).
Data Selection
@@
Data selection for fine-tuning has been extensively studied, focusing on improving quality (Li et al., 2023d), diversity (Liu et al., 2024), and distribution matching (Li et al., 2023e; Xia et al., 2024; Ni et al., 2024). For pretraining, various lightweight filters are utilized (Albalak et al., 2024), including heuristic-based (e.g., language and item count filtering), classifier-based (Brown et al., 2020), and perplexity-based approaches (Wenzek et al., 2019). The massive public RedPajama-Data-v2 dataset (Computer, 2023), for example, leverages over 40 quality indicators for data filtering and reweighting. Nevertheless, strict filtering like blocklist (Raffel et al., 2020) and Safety API filtering (Welbl et al., 2021), have been found to hurt evaluation loss or induce bias (Dodge et al., 2021). To our knowledge, we are the first to explore token-level data selection, aimed at enhancing data quality and information density at the most fundamental granularity.
Language Model Training Dynamics
@@
5 Discussion and Future Work
Generalization
In math continual pretraining, as depicted in Figure 6, training exclusively with SLM leads to quickly convergence to the domain focused by the reference model, accompanied by a significant rise in the loss of unselected tokens. Although no adverse effects, like biases, have been observed from the increased loss yet, a general pretraining loss on text and code may prevent overfitting (Goodhart and Goodhart, 1984), as suggested by Ouyang et al. (2022) and Azerbayev et al. (2023). Furthermore, future efforts could broaden the corpus scope of the reference model, and enlarge the pretraining data size, as exemplified by DeepSpeedMath (Shao et al., 2024a).
@@ 5 Discussion and Future Work
SLM may be extended to supervised fine-tuning to address the noise and distribution mismatches in many SFT datasets. Another potential application is alignment, e.g., by training a reference model to emphasize helpfulness, truthfulness, and harmlessness, we may obtain a base model that is natively aligned during the pretraining stage.
Appendix A Author Contributions
Appendix B Analysis and Visualization of Tokens in Pretraining
B.1 More Details of Four Categories Tokens
@@
Next, we can classify the tokens based on $\Delta\mathcal{L}$ and the $\mathcal{L}_{\text{mean}}$. We categorize tokens with $\Delta\mathcal{L}<-0.2$ as H→L (loss decreases from high to low) category tokens, and tokens with $\Delta\mathcal{L}>0.2$ as L→H (loss increases from low to high) category tokens. If $-0.2\leq\Delta\mathcal{L}\leq 0.2$ and $l_{n}\leq\mathcal{L}_{\text{mean}}$, then tokens are classified as L→L (loss remains low); if $l_{n}>\mathcal{L}_{\text{mean}}$, they are classified as H→H (loss remains high). In Figure 10, we visualize examples of the four categories of tokens in actual text.
B.2 Non-Converging Tokens in Pretrainig
@@
Appendix C Evalution Details
C.1 Math Evalution
@@
C.2 General Evalution
In the evaluation of general domain, we followed the lm-evaluation-harness (Gao et al., 2023) and evalute model on MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), AGIEval (Zhong et al., 2023), ARC-Easy and ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), OpenBookQA (Mihaylov et al., 2018). On HumanEval (Zheng et al., 2023) and TydiQA (Clark et al., 2020), we follow the evaluation pipeline of open-instrcut (Ivison et al., 2023) and report Pass@1 and Pass@10 for HumanEval and F1 for TydiQA. For MBPP (Austin et al., 2021) benchmark, we follow the evaluation pipeline of DeepSeek-Coder (Guo et al., 2024), and report Pass@1 and Pass@10.
Appendix D Relate the Selected Tokens’ Loss to Downstream Task Performance
@@
Appendix E Examples of Tokens Selected by SLM
E.1 Token Selected Examples
Figure 12: Specific examples of selecting tokens during the ToT training process of the Rho-1. The tokens marked in blue represent the actual tokens trained during the ToT training process, while the remaining black tokens are not trained during the ToT training process.
@@ E.1 Token Selected Examples
E.2 Dynamic Token Selected

### v2

@@ B.1 Pretraining Data Optimization
The objective of optimizing pre-training corpora is to maximize the performance and efficiency of language model training by improving the quality and scale of the pretrain data mixture. This includes data collecting through crawling (Raffel et al., 2020) or synthesis (Polu and Sutskever, 2020; Gunasekar et al., 2023), de-duplication (Lee et al., 2021; Kandpal et al., 2022; Tirumala et al., 2023), filtering and selection (Xie et al., 2024a; Albalak et al., 2024), as well as data composition (Xie et al., 2024b) and curriculum (Chen et al., 2024; Ma et al., 2024).
B.2 Data Selection
@@
Data selection for fine-tuning has been extensively studied, focusing on improving quality (Li et al., 2023c), diversity (Liu et al., 2024), and distribution matching (Li et al., 2023d; Xia et al., 2024; Ni et al., 2024; Kang et al., 2024). For pretraining, various lightweight filters are utilized (Albalak et al., 2024), including heuristic-based (e.g., language and item count filtering), classifier-based (Brown et al., 2020), and perplexity or loss-based approaches (Wenzek et al., 2019; Qin et al., 2024). The massive public RedPajama-Data-v2 dataset (Computer, 2023), for example, leverages over 40 quality indicators for data filtering and reweighting. Nevertheless, strict filtering like blocklist (Raffel et al., 2020) and Safety API filtering (Welbl et al., 2021), have been found to hurt evaluation loss or induce bias (Dodge et al., 2021).
Sample-level selection has been extensively studied in previous research, particularly through online batch selection methods (Loshchilov and Hutter, 2015; Schaul et al., 2015; Katharopoulos and Fleuret, 2018; Jiang et al., 2019). These approaches have been successfully applied to diverse classification tasks (Mindermann et al., 2022) and language modeling (Fan and Jaggi, 2023).
Token-level training strategies have also been explored, especially for the pre-training of BERT-like models using Masked Language Modeling (MLM) (Devlin et al., 2018). Specifically, “selective masking” involves masking important tokens in the input to focus on learning tokens that are more relevant to downstream tasks (Gu et al., 2020; Lad et al., 2022), whereas “token dropping” aims to reduce training costs by omitting less important tokens (Zhong et al., 2023a; Hou et al., 2022). Additionally, some research has approached the analysis and detection of under-trained tokens from a tokenization perspective (Rumbelow and Watkins, 2023; Land and Bartolo, 2024). To our knowledge, we are the first to explore token-level data selection for large language model training, aimed at enhancing data quality and information density at the most fundamental granularity.
B.3 Language Model Training Dynamics
@@ B.4 Scaling Laws
Appendix C Limitations and Future Work
Generalizability
In math continual pretraining, as depicted in Figure 6, training exclusively with SLM leads to quickly convergence to the domain focused by the reference model, accompanied by a significant rise in the loss of unselected tokens. Although no adverse effects, like biases, have been observed from the increased loss yet, a general pretraining loss on text and code may prevent overfitting (Goodhart and Goodhart, 1984), as suggested by Ouyang et al. (2022) and Azerbayev et al. (2023). Furthermore, future efforts could broaden the corpus scope of the reference model, and enlarge the pretraining data size, as exemplified by DeepSpeedMath (Shao et al., 2024).
@@ Appendix C Limitations and Future Work
SLM may be extended to supervised fine-tuning to address the noise and distribution mismatches in many SFT datasets. Another potential application is alignment, e.g., by training a reference model to emphasize helpfulness, truthfulness, and harmlessness, we may obtain a base model that is natively aligned during the pretraining stage. Meanwhile, we believe that the idea of SLM may find broader applications in multimodal data such as images, videos, and speech, which have a high noise-to-information ratio than text.
Appendix D Analysis and Visualization of Tokens in Pretraining
Figure 10: The loss of four categories of tokens during Mistral-7B pretraining on OpenWebMath. (a) shows the loss of H→H, L→H, H→L, and L→L tokens during pretraining. (b) and (c) show three cases of fluctuating tokens’ loss in L→L and H→H during pretraining, respectively.
D.1 More Details of Four Categories Tokens
@@
Next, we can classify the tokens based on $\Delta\mathcal{L}$ and the $\mathcal{L}_{\text{mean}}$. We categorize tokens with $\Delta\mathcal{L}<-0.2$ as H→L (loss decreases from high to low) category tokens, and tokens with $\Delta\mathcal{L}>0.2$ as L→H (loss increases from low to high) category tokens. If $-0.2\leq\Delta\mathcal{L}\leq 0.2$ and $l_{n}\leq\mathcal{L}_{\text{mean}}$, then tokens are classified as L→L (loss remains low); if $l_{n}>\mathcal{L}_{\text{mean}}$, they are classified as H→H (loss remains high). In Figure 10, we have added the tokens’ loss curves of the 7B model which is consistent with the other experimental settings in §2.1, for readers to refer to whether similar phenomena exist on larger models. In Figure 11, we visualize examples of the four categories of tokens in actual text.
D.2 Non-Converging Tokens in Pretrainig
@@
Appendix E Evalution Details
E.1 Math Evalution
@@
E.2 General Evalution
In the evaluation of general domain, we followed the lm-evaluation-harness (Gao et al., 2023) and evalute model on MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), AGIEval (Zhong et al., 2023b), ARC-Easy and ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), OpenBookQA (Mihaylov et al., 2018). On HumanEval (Zheng et al., 2023) and TydiQA (Clark et al., 2020), we follow the evaluation pipeline of open-instrcut (Ivison et al., 2023) and report Pass@1 and Pass@10 for HumanEval and F1 for TydiQA. For MBPP (Austin et al., 2021) benchmark, we follow the evaluation pipeline of DeepSeek-Coder (Guo et al., 2024), and report Pass@1 and Pass@10.
Appendix F Relate the Selected Tokens’ Loss to Downstream Task Performance
@@
Appendix G Examples of Tokens Selected by SLM
G.1 Token Selected Examples
@@
G.2 Dynamic Token Selected
Appendix H Self-Reference Setting
Table 4: Full Self-Reference results on Tinyllama-1.1B.
Score
Function
Select
Ratio
GSM8K MATH SVAMP ASDiv MAWPS MQA AVG
- 100% 6.3 2.6 21.7 36.7 47.7 13.9 21.5
$\mathcal{L}_{\Delta}(x_{i})$ 90% 7.4 4.4 23.4 38.7 51.9 14.4 23.4
80% 6.4 4.6 23.1 39.7 52.0 14.3 23.4
70% 6.7 4.6 23.3 40.0 54.5 14.3 23.9
60% 7.0 4.6 22.2 38.5 52.2 13.7 23.0
50% 5.7 4.2 20.7 36.7 46.7 10.3 20.7
$\mathcal{H}_{\text{RM}}(x_{i})$ 90% 6.7 3.0 23.7 40.3 52.3 13.1 23.2
80% 6.8 3.6 22.5 40.6 52.9 13.6 23.3
70% 7.0 4.8 23.0 39.3 50.5 13.5 23.0
60% 6.5 4.8 26.5 37.3 49.7 15.6 23.4
50% 4.7 5.8 20.9 33.8 42.5 11.1 19.8
$\mathcal{H}_{\text{RM}}(x_{i})\cup\mathcal{L}_{\Delta}(x_{i})$ $50\%\cup 70\%(80\%)$ 6.4 3.6 22.7 38.4 52.6 15.3 23.2
$70\%\cup 60\%(77\%)$ 6.3 4.6 24.4 39.6 51.4 16.3 23.8
$70\%\cup 50\%(75\%)$ 6.9 5.6 23.2 39.9 52.9 12.6 23.5
$60\%\cup 60\%(70\%)$ 6.7 5.2 24.7 39.2 50.6 14.6 23.5
$60\%\cup 50\%(68\%)$ 7.1 5.8 21.7 37.3 49.6 15.3 22.8
$60\%\cup 40\%(65\%)$ 7.3 6.0 23.6 36.9 48.6 13.1 22.6
$\mathcal{H}_{\text{RM}}(x_{i})\cap\mathcal{L}_{\Delta}(x_{i})$ $80\%\cap 90\%(76\%)$ 6.0 4.4 23.7 38.5 51.2 13.3 22.8
$75\%\cap 75\%(72\%)$ 7.8 5.2 24.2 39.4 54.9 14.7 24.4
$70\%\cap 90\%(68\%)$ 6.8 4.6 22.2 40.3 53.0 14.8 23.6
$80\%\cap 80\%(67\%)$ 8.2 6.4 21.2 39.1 53.4 15.0 23.9
$70\%\cap 70\%(60\%)$ 7.1 5.0 23.5 41.2 53.8 18.0 24.8
In this section, we will provide a detailed introduction to the reference loss score function and information entropy score function in SLM. Reference loss score function is to directly use the loss of the reference model as the basis for selecting tokens. The higher the token’s loss of the reference model, the lower the expectation that the token will be selected. The score $\mathcal{L}_{\text{RM}}(x_{i})$ can be directly obtained by referring to Equation 1. Information entropy score function is to select the corresponding token based on the information entropy of the reference model in each token. The information entropy of token $x_{i}$ can be expressed as:
$\mathcal{H}_{\text{RM}}(x_{i})=-\sum_{k=1}^{V}P(t_{k}|x_{<i})\log P(t_{k}|x_{<i}),$ (8)
where $t_{k}$ represents the i-th token in the vocabulary, and $V$ represents the size of the vocabulary. The intuition of this strategy is that the higher the information entropy, the higher the uncertainty of the token in the context. Therefore, we consider that if the language model is still uncertain for certain tokens after pretraining, we do not hope that the language model to learn it during pretraining. In Table 4, we provide more SLM results, including different select ratios and combinations of two score functions, for the convenience of readers to refer to.
Appendix I Weak-to-Strong Generalization
Table 5: Weak-to-Strong generalization result on math benchmark.
Model Train Toks GSM8K MATH SVAMP ASDiv MAWPS TAB MQA
MMLU
STEM
SAT AVG
Llama-2-7B-CT 15B 28.4 13.6 50.3 62.8 79.5 37.6 34.1 41.6 43.5 43.5
Llama-2-7B-CT w/ 1B RM 10.5B 29.8 16.0 55.5 63.7 80.4 37.9 34.3 38.2 43.8 44.4
Apart from the main experiments where we use the same base model for the reference and continual pretraining, we also investigate if a smaller reference model can effectively guide the pretraining of a larger model. We use Tinyllma-1.1B as reference model and continual pretraining Llama-2-7B on 15B OpenWebMath tokens. Results presented in Table 5 indicate that, despite the considerable gap between the small and large models (Li et al., 2023e), employing the small reference model to token selection can still yield benefits to the pre-training of the larger model. If reference and training models have different vocabularies, one can consider performing token alignment (Wan et al., 2024; Fu et al., 2023), which we leave for future work.
Figure 13: Specific examples of selecting tokens during the selective pretraining process of the Rho-1. The tokens marked in blue represent the actual tokens trained during the training process, while the remaining black tokens are not trained during the training process.